# Gamma-ray spectra of the Crab, Vela and Geminga pulsars fitted with SED of the emission from their current sheet

### 24 Jan 2024 v2

@@
11institutetext: Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK
email: ardavan@ast.cam.ac.uk
We show that the spectral energy distribution (SED) of the tightly focused radiation generated by the superluminally moving current sheet in the magnetosphere of a non-aligned neutron star fits the gamma-ray spectra of the Crab, Vela and Geminga pulsars over the entire range of photon energies so far detected by Fermi-LAT, MAGIC and H.E.S.S. from them: over $10^{2}$ MeV to $20$ TeV. While emblematic of any emission that entails caustics, the SED introduced here radically differs from those of the disparate emission mechanisms currently invoked in the literature to fit the data in different sections of these spectra. We specify, moreover, the connection between the values of the fit parameters for the analysed spectra and the physical characteristics of the central neutron stars of the Crab, Vela and Geminga pulsars and their magnetospheres.
Gamma-ray spectra of the Crab, Vela and Geminga pulsars
fitted with SED of the emission from their current sheet
Houshang Ardavan
Key Words.:
pulsars: individual: J0534+2200, J0835-4510, J0633+1746 – gamma-rays: stars – stars: neutron – methods: data analysis – radiation mechanisms: non-thermal
@@ 1 Introduction
It is now firmly established that the magnetosphere of a non-aligned neutron star contains a rigidly rotating current sheet outside its light cylinder (see the review article Philippov & Kramer 2022 and the references therein). That this current sheet moves faster than light in vacuum is not incompatible with the requirements of special relativity because its superluminally moving distribution pattern is created by the correlated motion of subluminally moving charged particles (Bolotovskii & Bykov, 1990). The field generated by a constituent volume element of this current sheet embraces a synergy between the superluminal version of the field of synchrotron radiation and the vacuum version of the field of Čerenkov radiation (Ardavan, 2023, Sect. 2): the waves emanating from it possess a two-sheeted cusped envelope that spirals outward around the light cylinder all the way into the far zone (Ardavan, 2023, Figs. 3a and 4).
Only a single wave front propagates past an observer outside the envelope at any given time. But an observer inside the envelope receives three intersecting wave fronts, emanating from three distinct retarded positions of the source element simultaneously (Ardavan, 2023, Figs. 3b and 4). On each sheet of the envelope, two of these wave fronts coalesce and interfere constructively to generate an infinitely large radiation field in the time domain (Ardavan, 2021, Sect. 3.4). On the cusp locus of the envelope, where all three of these wave fronts (together with the three corresponding images of the source element) coalesce, the resulting radiation field has even a higher-order singularity (Ardavan, 2023, Figs. 5 and 6). Divergence of the field of a source point on the envelope and its cusp, which stems from the relativistic restrictions inherent in electrodynamics, reflects the fact that no superluminal source can be point-like. Superposition of the contributions from the constituent volume elements of the current sheet results in a field that, while sharply peaked, is finite (Ardavan, 2021, Sect. 4).
The field of the entire volume of the source receives its main contributions from those volume elements that approach the observer along the radiation direction with the speed of light and zero acceleration at the retarded time (Ardavan, 2021, Sect. 4.2). Phases of the waves emitted by such volume elements of the source possess two nearby stationary points that draw closer to each other the farther the waves are from their source (Ardavan, 2023, Fig. 7) and eventually coalesce at infinity. As a result, the waves in question constitute progressively more focused pulses, which by virtue of having time-domain profiles whose peaks get narrower the farther they propagate away from their source, embody the high-frequency component of the radiation from the current sheet (Ardavan, 2021, Sect. 5). The lower-frequency emission in the radio loud gamma-ray pulsars arises from the isolated stationary points of the phases of these waves (Ardavan, 2021, Sect. 5.4).
@@
A detailed analysis of the radiation that is generated by the superluminally moving current sheet in the magnetosphere of a non-aligned neutron star can be found in Ardavan (2021). A heuristic account of the mathematical results of that analysis in more transparent physical terms is presented in Ardavan (2022) and Ardavan (2023, Sect. 2).
@@
We begin here by deriving the spectral energy distribution (SED) of the most tightly focused component of the radiation that is emitted by the magnetospheric current sheet (Sect. 2). We then specify the values of the free parameters in the derived expression for which this SED best fits the data on the observed gamma-ray spectra of the Crab, Vela and Geminga pulsars (Sect. 3). The specified values of the fit parameters will be used, in conjunction with the results of the analysis in Ardavan (2021), to determine certain attributes of the central neutron stars of these pulsars and their magnetospheres in Sect. 4. The radical departure of the single explanation given here for the entire breadths of the analysed spectra from those normally given in terms of disjointed sets of ad hoc spectral distribution functions (such as simple or broken power-law functions with exponential cutoffs; see, e.g. Zanin 2017) and disparate emission mechanisms (such as synchro-curvature processes, inverse Compton scattering, or magnetic reconnection; see, e.g. The H.E.S.S. Collaboration et al. 2023) is briefly discussed in the final section of the paper (Sect. 5).
@@ 2 SED of the caustics generated by the superluminally moving current sheet
The frequency spectrum of the radiation that is generated by the superluminally moving current sheet in the magnetosphere of a non-aligned neutron star was presented, in its general form, in Ardavan (2021, Sect. 5.3). Here we derive the SED of the most tightly-focused component of this radiation from the general expression given in Eq. (177) of that paper.
In a case where the magnitudes of the vectors denoted by $\boldsymbol{\cal P}_{l}$ and $\boldsymbol{\cal Q}_{l}$ in Eq. (177) of Ardavan (2021) are appreciably larger than those of their counterparts, ${\bar{\boldsymbol{\cal P}}}_{l}$ and ${\bar{\boldsymbol{\cal Q}}}_{l}$, and the dominant contribution towards the Poynting flux $S_{\nu}$ of the radiation at the observation point is made by the term corresponding to $l=2$, Eq. (177) of Ardavan (2021) can be written as
@@
where Ai and ${\rm Ai}^{\prime}$ are the Airy function and the derivative of the Airy function with respect to its argument, respectively, $k=2\pi\nu/\omega$ is the frequency $\nu$ of the radiation in units of the rotation frequency $\omega/2\pi$ of the central neutron star, and $\kappa_{0}$ and $\sigma_{21}$ are two positive scalars. In the high-frequency regime $k\gg 1$, the coefficients of the Airy functions in the above expression stand for the complex vectors $\boldsymbol{\cal P}_{2}=k^{-1/2}\boldsymbol{\cal P}_{2}^{(2)}$ and $\boldsymbol{\cal Q}_{2}=k^{-1/2}\boldsymbol{\cal Q}_{2}^{(2)}$ in which $\boldsymbol{\cal P}_{2}^{(2)}$ and $\boldsymbol{\cal Q}_{2}^{(2)}$ are defined by Eqs (138)-(146) of Ardavan (2021). The variable $\sigma_{21}$ has a vanishingly small value at those privileged colatitudes (relative to the star’s spin axis) where the high-frequency radiation is observable (Ardavan, 2021, Sect. 4.5). For the purposes of the analysis in this paper, we may therefore replace $\boldsymbol{\cal P}_{2}^{(2)}$ and $\boldsymbol{\cal Q}_{2}^{(2)}$ by their limiting values for $k\gg 1$ and $0\leq\sigma_{21}\ll 1$ and treat them as constant parameters.
@@
$\nu S_{\nu}$ $=$ $\frac{\omega\kappa_{0}}{2\pi}k^{-2/3}\left|\boldsymbol{\cal P}^{(2)}_{2}\,{\rm Ai}(-k^{2/3}\sigma_{21}^{2})\right.$ (2)
$\left.-{\rm i}k^{-1/3}\boldsymbol{\cal Q}^{(2)}_{2}\,{\rm Ai}^{\prime}(-k^{2/3}\sigma_{21}^{2})\right|^{2}.$
@@
$\kappa_{1}=\frac{\omega\kappa_{0}}{2\pi}\left|\boldsymbol{\cal P}^{(2)}_{2}\right|^{2},\,\,\,\zeta_{1}=\frac{\left|{\boldsymbol{\cal Q}^{(2)}_{2}}\right|}{\left|{\boldsymbol{\cal P}^{(2)}_{2}}\right|},\,\,\,\cos\beta=\frac{\Im\left(\boldsymbol{\cal Q}^{(2)}_{2}\cdot\boldsymbol{\cal P}^{(2)*}_{2}\right)}{\left|\boldsymbol{\cal Q}^{(2)}_{2}\right|\left|\boldsymbol{\cal P}^{(2)}_{2}\right|},$ (4)
and $\Im{}$ and *** denote an imaginary part and the complex conjugate, respectively. The above spectrum is emblematic of any radiation that entails caustics (see Stamnes, 1986).
To take account of the fact that the parameter $\sigma_{21}$ assumes a non-zero range of values across the (non-zero) latitudinal width of the detected radiation beam (Ardavan, 2021, Sect. 4.5), we must integrate $\nu S_{\nu}$ with respect to $\sigma_{21}$ over a finite interval $\rho\sigma_{0}\leq\sigma_{21}\leq\sigma_{0}$ with $\sigma_{0}\ll 1$ and $10\leq\rho<$ 1. Performing the integration of the Airy functions in Eq. (3) with respect to $\sigma_{21}$ by means of Mathematica, we thus obtain
${\cal S}_{\nu}$ $=$ $\nu\int_{\rho\sigma_{0}}^{\sigma_{0}}S_{\nu}\,{\rm d}\sigma_{21}$ (5)
$=$ $\kappa\left[f_{1}(\chi,\rho)+\frac{\zeta^{2}}{4\sqrt{3}}f_{2}(\chi,\rho)-\frac{\zeta\cos\beta}{2\sqrt{3}}f_{3}(\chi,\rho)\right],$
@@
$f_{1}$ $=$ $\bigg{[}3\Gamma\left(\frac{7}{6}\right)\eta\chi^{-2/3}{}_{2}F_{3}\left(\begin{matrix}1/6&1/6&{}\\ 1/3&2/3&7/6\end{matrix};-\eta^{6}\chi^{2}\right)$ (15)
$+\pi^{1/2}\eta^{3}{}_{2}F_{3}\left(\begin{matrix}1/2&1/2&{}\\ 2/3&4/3&3/2\end{matrix}\,;-\eta^{6}\chi^{2}\right)$
@@
and ${}_{2}F_{3}$ and ${}_{24}G^{31}$ are respectively the generalised hypergeometric function (see Olver et al., 2010) and the generalised Meijer G-Function 111 https://mathworld.wolfram.com/MeijerG-Function.html. The variable $\chi$ that appears in the above expressions is related to the frequency $\nu$ of the radiation via $\chi=4\pi{\sigma_{0}}^{3}\nu/(3\omega)$.
The scale and shape of the SED given in Eq. (5) depend on the five parameters $\kappa$, $\zeta$, $\beta$, $\sigma_{0}$ and $\rho$: parameters whose values are dictated by the characteristics of the magnetospheric current sheet (see Sect. 4). The parameters $\zeta$, $\beta$ and $\rho$ determine the shape of the energy distribution while the parameters $\kappa$ and $\sigma_{0}$ determine the position of this distribution along the energy-flux and photon-energy axes, respectively.
Figure 1: Phase-averaged gamma-ray spectrum of the Crab pulsar. The FERMI-LAT data points below $25$ GeV (coloured black) are those reported in Abdo et al. (2010). The MAGIC data points between $25$ and $100$ GeV (coloured blue) are those reported in Aleksić et al. (2011). The MAGIC data points beyond $100$ Gev (coloured green) are extracted, by the procedure described in Ardavan (2023, Sect. 4), from those reported by Ansoldi et al. (2016). The curve (coloured red) is a plot of the SED described by Eq. (5) for the parameters given in Eq. (19).
@@ 3 Fits to the gamma-ray data on the SEDs of the Crab, Vela and Geminga pulsars
In this section we use Mathematica’s ‘NonlinearModelFit’ procedure222 https://reference.wolfram.com/language/ref/NonlinearModelFit.html and the statistical information that it provides to determine the values of the fit parameters and their standard errors. Where, owing to the complexity of the expression in Eq. (5), this procedure breaks down and so the fits to the data are obtained by trial and error, only the values of these parameters are specified. The fit residuals in the case of each pulsar turn out to be smaller than the corresponding observational errors for almost all data points (see Figs. 13). Since there are no two values of any of the fit parameters for which the present SED has the same shape and position, the specified values of the fit parameters are moreover unique.
@@ 3.1 Spectrum of the Crab pulsar (PSR J0534+2200)
Figure 1 shows the data on the energy spectrum of the phase-averaged gamma-ray emission from the Crab pulsar (Abdo et al., 2010; Aleksić et al., 2011; Ansoldi et al., 2016) and a plot of the SED described by Eq. (5) that fits them best. The parameters for which the function ${\cal S}_{\nu}(\chi)$ (depicted by the red curve in this figure) is plotted have the following values and standard errors:
@@
$\sigma_{0}$ $=$ $(4.76\pm 2.61)\times 10^{-8},\,\,\,\zeta=0.394\pm 0.076,$
@@
where $h$ stands for the Planck constant and $2\pi/\omega$ has been set equal to the period of the Crab pulsar ($0.033$ s). Figure 1 was first reported in Ardavan (2023) but with an incorrect label on its vertical axis.
As the photon energy $h\nu$ increases beyond $1.2$ TeV, the slope of the curve in Fig. 1 continues to decrease at the relatively slow rate in which it decreases past $50$ GeV. Thus the upper limits given by Ansoldi et al. (2016) for the flux density of gamma-rays whose energies exceed $1.2$ TeV all lie above the continuation of the curve shown in Fig. 1.
Figure 2: Gamma-ray spectrum of the Vela pulsar. The FERMI-LAT data points below $10^{2}$ GeV (coloured black) are those reported in Abdo et al. (2013). The H.E.S.S. data above $260$ GeV (shown as the magenta area between $1\sigma$ uncertainty bands and the magenta upper limits) are those reported in The H.E.S.S. Collaboration et al. (2023). The curve (coloured red) is a plot of the SED described by Eq. (5) for the parameters given in Eqs. (20) and (21).

### 03 Mar 2024 v3

@@
Abstract
We show that the spectral energy distribution (SED) of the tightly focused radiation generated by the superluminally moving current sheet in the magnetosphere of a non-aligned neutron star fits the gamma-ray spectra of the Crab, Vela and Geminga pulsars over the entire range of photon energies so far detected by Fermi-LAT, MAGIC and HESS from them: over $10^{2}$ MeV to $20$ TeV. While emblematic of any emission that entails caustics, the SED introduced here radically differs from those of the disparate emission mechanisms currently invoked in the literature to fit the data in different sections of these spectra. We specify, moreover, the connection between the values of the fit parameters for the analysed spectra and the physical characteristics of the central neutron stars of the Crab, Vela and Geminga pulsars and their magnetospheres.
Keywords— pulsars: individual: J0534+2200, J0835-4510, J0633+1746 – gamma-rays: stars – stars: neutron – methods: data analysis – radiation mechanisms: non-thermal
@@ 1 Introduction
Attempts at explaining the radiation from pulsars has so far been focused mainly on mechanisms of acceleration of charged particles (see, e.g. the references in [1] and [2]): an approach spurred by the fact that, once the relevant version of this mechanism is identified, one can calculate the electric current density associated with the accelerating charged particles involved and thereby evaluate the classical expression for the retarded potential that describes the looked-for radiation. In the present paper, however, we evaluate the retarded potential, and hence the generated radiation field, using the macroscopic distribution of electric charge-current density that is already provided by the numerical computations of the structure of a non-aligned pulsar magnetosphere (section 2 of [3]). Both the radiation field thus calculated and the electric and magnetic fields that pervade the pulsar magnetosphere are solutions of Maxwell’s equations for the same charge-current distribution. These two solutions are completely different, nevertheless, because they satisfy different boundary conditions: the far-field boundary conditions with which the structure of the pulsar magnetosphere is computed are radically different from the corresponding boundary conditions with which the retarded solution of these equations (i.e. the solution describing the radiation from the charges and currents in the pulsar magnetosphere) is derived (see section 3 and the last paragraph in section 6 of [3]).
Numerical computations based on the force-free and particle-in-cell formalisms have now firmly established that the magnetosphere of a non-aligned neutron star entails a current sheet outside its light cylinder whose rotating distribution pattern moves with linear speeds exceeding the speed of light in vacuum (see [4, 5, 6] and the references in [7]). However, the role played by the superluminal motion of this current sheet in generating the multi-wavelength, focused pulses of radiation that we receive from neutron stars is not generally acknowledged. Given that the superluminally moving distribution pattern of this current sheet is created by the coordinated motion of aggregates of subluminally moving charged particles [8, 9, 10], the motion of any of its constituent particles is too complicated to be taken into account individually. Only the densities of charges and currents enter Maxwell’s equations, on the other hand, so that the macroscopic charge-current distribution associated with the magnetospheric current sheet takes full account of the contributions toward the radiation that arise from the complicated motions of the charged particles comprising it.
The radiation field generated by a uniformly rotating volume element of the distribution pattern of the current sheet in the magnetosphere of a non-aligned neutron star embraces a synergy between the superluminal version of the field of synchrotron radiation and the vacuum version of the field of Čerenkov radiation. Once superposed to yield the emission from the entire volume of the source, the contributions from the volume elements of this distribution pattern that approach the observation point with the speed of light and zero acceleration at the retarded time interfere constructively and form caustics in certain latitudinal directions relative to the spin axis of the neutron star. The waves that embody these caustics are more focused the farther they are from their source: as their distance from their source increases, two nearby stationary points of their phases draw closer to each other and eventually coalesce at infinity. By virtue of their narrow peaks in the time domain, the resulting focused pulses thus procure frequency spectra whose distributions extend from radio waves to gamma-rays [3, 11, 12].
@@
A detailed analysis of the radiation that is generated by the superluminally moving current sheet in the magnetosphere of a non-aligned neutron star can be found in [3]. Heuristic accounts of the mathematical results of that analysis in more transparent physical terms are presented in [11] and section 2 of [12].
@@
We begin here by deriving the spectral energy distribution (SED) of the most tightly focused component of the radiation that is emitted by the magnetospheric current sheet (section 2). We then specify the values of the free parameters in the derived expression for which this SED best fits the data on the observed gamma-ray spectra of the Crab, Vela and Geminga pulsars (section 3). The specified values of the fit parameters will be used, in conjunction with the results of the analysis in [3], to determine certain attributes of the central neutron stars of these pulsars and their magnetospheres in section 4. The radical departure of the single explanation given here for the entire breadths of the analysed spectra from those normally given in terms of disjointed sets of ad hoc spectral distribution functions (such as simple or broken power-law functions with exponential cutoffs; see, e.g. [13]) and disparate emission mechanisms (such as synchro-curvature processes, inverse Compton scattering, or magnetic reconnection; see, e.g. [2]) is briefly discussed in the final section of the paper (section 5).
@@ 2 SED of the caustics generated by the superluminally moving current sheet
The frequency spectrum of the radiation that is generated by the superluminally moving current sheet in the magnetosphere of a non-aligned neutron star was presented, in its general form, in section 5.3 of [3]. Here we derive the SED of the most tightly-focused component of this radiation from the general expression given in eq. (177) of that paper.
In a case where the magnitudes of the vectors denoted by $\boldsymbol{\cal P}_{l}$ and $\boldsymbol{\cal Q}_{l}$ in eq. (177) of [3] are appreciably larger than those of their counterparts, ${\bar{\boldsymbol{\cal P}}}_{l}$ and ${\bar{\boldsymbol{\cal Q}}}_{l}$, and the dominant contribution towards the Poynting flux $S_{\nu}$ of the radiation at the frequency $\nu$ is made by only one of the two terms corresponding to $l=1$ and $l=2$, e.g. $l=2$, eq. (177) of [3] can be written as
@@
where Ai and ${\rm Ai}^{\prime}$ are the Airy function and the derivative of the Airy function with respect to its argument, respectively, $k=2\pi\nu/\omega$ is the frequency $\nu$ of the radiation in units of the rotation frequency $\omega/2\pi$ of the central neutron star, and $\kappa_{0}$ and $\sigma_{21}$ are two positive scalars. In the high-frequency regime $k\gg 1$, the coefficients of the Airy functions in the above expression stand for the complex vectors $\boldsymbol{\cal P}_{2}=k^{-1/2}\boldsymbol{\cal P}_{2}^{(2)}$ and $\boldsymbol{\cal Q}_{2}=k^{-1/2}\boldsymbol{\cal Q}_{2}^{(2)}$ in which $\boldsymbol{\cal P}_{2}^{(2)}$ and $\boldsymbol{\cal Q}_{2}^{(2)}$ are defined by eqs. (138)-(146) of [3]. The variable $\sigma_{21}$ has a vanishingly small value at those privileged colatitudes (relative to the star’s spin axis) where the high-frequency radiation is observable (section 4.5 of [3]). For the purposes of the analysis in this paper, we may therefore replace $\boldsymbol{\cal P}_{2}^{(2)}$ and $\boldsymbol{\cal Q}_{2}^{(2)}$ by their limiting values for $k\gg 1$ and $0\leq\sigma_{21}\ll 1$ and treat them as constant parameters.
@@
$\nu S_{\nu}=\frac{\omega\kappa_{0}}{2\pi}k^{-2/3}\left|\boldsymbol{\cal P}^{(2)}_{2}\,{\rm Ai}(-k^{2/3}\sigma_{21}^{2})-{\rm i}k^{-1/3}\boldsymbol{\cal Q}^{(2)}_{2}\,{\rm Ai}^{\prime}(-k^{2/3}\sigma_{21}^{2})\right|^{2}.$ (2)
@@
$\kappa_{1}=\frac{\omega\kappa_{0}}{2\pi}\left|\boldsymbol{\cal P}^{(2)}_{2}\right|^{2},\quad\zeta_{1}=\frac{\left|{\boldsymbol{\cal Q}^{(2)}_{2}}\right|}{\left|{\boldsymbol{\cal P}^{(2)}_{2}}\right|},\quad\cos\beta=\frac{\Im\left(\boldsymbol{\cal Q}^{(2)}_{2}\cdot\boldsymbol{\cal P}^{(2)*}_{2}\right)}{\left|\boldsymbol{\cal Q}^{(2)}_{2}\right|\left|\boldsymbol{\cal P}^{(2)}_{2}\right|},$ (4)
and $\Im{}$ and *** denote an imaginary part and the complex conjugate, respectively. The above spectrum is emblematic of any radiation that entails caustics (see [14]).
To take account of the fact that the parameter $\sigma_{21}$ assumes a non-zero range of values across the (non-zero) latitudinal width of the detected radiation beam (section 4.5 of [3]), we must integrate $\nu S_{\nu}$ with respect to $\sigma_{21}$ over a finite interval $\rho\sigma_{0}\leq\sigma_{21}\leq\sigma_{0}$ with $\sigma_{0}\ll 1$ and $10\leq\rho<$ 1. Performing the integration of the Airy functions in eq. (3) with respect to $\sigma_{21}$ by means of Mathematica, we thus obtain
${\cal S}_{\nu}=\nu\int_{\rho\sigma_{0}}^{\sigma_{0}}S_{\nu}\,{\rm d}\sigma_{21}=\kappa\left[f_{1}(\chi,\rho)+\frac{\zeta^{2}}{4\sqrt{3}}f_{2}(\chi,\rho)-\frac{\zeta\cos\beta}{2\sqrt{3}}f_{3}(\chi,\rho)\right],$ (5)
@@
$f_{1}$ $=$ $\bigg{[}3\Gamma\left(\frac{7}{6}\right)\eta\chi^{-2/3}{}_{2}F_{3}\left(\begin{matrix}1/6&1/6&{}\\ 1/3&2/3&7/6\end{matrix};-\eta^{6}\chi^{2}\right)+\pi^{1/2}\eta^{3}{}_{2}F_{3}\left(\begin{matrix}1/2&1/2&{}\\ 2/3&4/3&3/2\end{matrix}\,;-\eta^{6}\chi^{2}\right)$ (14)
@@
and ${}_{2}F_{3}$ and ${}_{24}G^{31}$ are respectively the generalised hypergeometric function [15] and the generalised Meijer G-Function.222 https://mathworld.wolfram.com/Meijer G-Function.html The variable $\chi$ that appears in the above expressions is related to the frequency $\nu$ of the radiation via $\chi=4\pi{\sigma_{0}}^{3}\nu/(3\omega)$.
The scale and shape of the SED given in eq. (5) depend on the five parameters $\kappa$, $\zeta$, $\beta$, $\sigma_{0}$ and $\rho$: parameters whose values are dictated by the characteristics of the magnetospheric current sheet (see section 4). The parameters $\zeta$, $\beta$ and $\rho$ determine the shape of the energy distribution while the parameters $\kappa$ and $\sigma_{0}$ determine the position of this distribution along the energy-flux and photon-energy axes, respectively.
Figure 1: Phase-averaged gamma-ray spectrum of the Crab pulsar. The Fermi-LAT data points below $25$ GeV (coloured black) are those reported in [16]. The MAGIC data points between $25$ and $100$ GeV (coloured blue) are those reported in [17]. The MAGIC data points beyond $100$ Gev (coloured green) are extracted, by the procedure described in section 4 of [12], from those reported by [18]. The curve (coloured red) is a plot of the SED described by eq. (5) for the parameters given in eq. (18).
@@ 3 Fits to the gamma-ray data on the SEDs of the Crab, Vela and Geminga pulsars
In this section we use Mathematica’s ‘NonlinearModelFit’ procedure333 https://reference.wolfram.com/language/ref/NonlinearModelFit.html and the statistical information that it provides to determine the values of the fit parameters and their standard errors. Where, owing to the complexity of the expression in eq. (5), this procedure fails to work and so the fits to the data are obtained by elementary iteration, only the values of these parameters are specified. The fit residuals in the case of each pulsar turn out to be smaller than the corresponding observational errors for almost all data points (see figures 13). Since there are no two values of any of the fit parameters for which the present SED has the same shape and position, the specified values of the fit parameters are moreover unique.
@@ 3.1 Spectrum of the Crab pulsar (PSR J0534+2200)
Figure 1 shows the data on the energy spectrum of the phase-averaged gamma-ray emission from the Crab pulsar [16, 17, 18] and a plot of the SED described by eq. (5) that fits them best. The parameters for which the function ${\cal S}_{\nu}(\chi)$ (depicted by the red curve in this figure) is plotted have the following values and standard errors:
@@
$\sigma_{0}$ $=$ $(4.76\pm 0.28)\times 10^{-8},\,\,\,\zeta=0.394\pm 0.076,$
@@
in which $h$ and $(h\nu)_{\rm GeV}$ stand for the Planck constant and photon energy in units of GeV, respectively, and $2\pi/\omega$ has been set equal to the period of the Crab pulsar ($0.033$ s). (Figure 1 was first reported in [12] but with an incorrect label on its vertical axis.)
As the photon energy $h\nu$ increases beyond $1.2$ TeV, the slope of the curve in figure 1 continues to decrease at the relatively slow rate in which it decreases past $50$ GeV. Thus the upper limits given in [18] for the flux density of gamma-rays whose energies exceed $1.2$ TeV all lie above the continuation of the curve shown in figure 1.
Figure 2: Gamma-ray spectrum of the Vela pulsar. The Fermi-LAT data points below $10^{2}$ GeV (coloured black) are those reported in [19]. The HESS data above $260$ GeV (shown as the magenta area between $1\sigma$ uncertainty bands and the magenta upper limits) are those reported in [2]. The curve (coloured red) is a plot of the SED described by eq. (5) for the parameters given in eqs. (19) and (20).